# Large Language Model Enabled Multi-Task Physical Layer Network

### v1

@@ III-B Input Encoders
Adapting text-based pre-trained LLM to multiple communication domain tasks is another challenging problem. To be specific, firstly, there is a huge gap between specific characteristics of communication domain data and the natural language. Therefore, directly applying a general-domain LLM to these tasks may lead to poor performance. Secondly, the format, length and meaning of the data also differs significantly among the tasks. The difficulties urge us to design dedicated input encoder modules for each task to fit for the feature space of the LLM, and facilitate multi-task feature extraction.
@@ III-B1 Encoder for multi-user precoding
Since neural networks generally deal with real numbers, we first convert the complex multi-user channel into real tensors $\mathbf{X}_{\rm Pre}^{\rm in}\in\mathbb{R}^{K\times 2N_{t}}$ as the input. For multi-user precoding, we employ a transfomer to extract the channel feature and inter-user relationship. The transfomer is composed of $L=3$ blocks, the structure of which is presented in Fig. 2 , including a multi-head self-attention module and a multilayer perceptron (MLP) module. The input $\mathbf{X}_{\rm Pre}^{\rm in}$ is first processed by the multi-head self-attention module:
@@
which serves as the input of next block. Denote the output of the transfomer as $\mathbf{X}_{\rm Pre}^{\rm mlp(3)}$, and then a linear projection layer projects $\mathbf{X}_{\rm Pre}^{\rm mlp(3)}$ into the language model space
@@ III-B2 Encoder for signal detection
The input of signal detection contains the uplink channel $\bar{\mathbf{H}}^{m}$ and the received signal $\bar{\mathbf{y}}^{m}$. The encoder rearranges the channel $\bar{\mathbf{H}}^{m}$ as one token $\mathbf{X}_{\rm Dec}^{\rm Hin}\in\mathbb{R}^{1\times 2KN_{t}}$. Then a MLP module is employed to extract features from the channel
$\mathbf{X}_{\rm Dec}^{\rm Hmlp}={\rm MLP}(\mathbf{X}_{\rm Dec}^{\rm Hin}).$ (15)
We also extract shallow features from the rearranged real received signal $\mathbf{X}_{\rm Dec}^{\rm yin}$ by a MLP module
$\mathbf{X}_{\rm Dec}^{\rm ymlp}={\rm MLP}(\mathbf{X}_{\rm Dec}^{\rm yin}).$ (16)
@@
$\mathbf{X}_{\rm Dec}^{\rm emb}={\rm Linear}([\mathbf{X}_{\rm Dec}^{\rm Hmlp},\mathbf{X}_{\rm Dec}^{\rm ymlp}]).$ (17)
@@ III-B3 Encoder for channel prediction
Consider that the CSI $\mathbf{H}^{t}\in\mathbb{C}^{N_{t}\times M}$ of at time $t$ is the high-dimensional structural data, directly predicting the matrix by the network will bring significant complexity and training time, especially for a large number of antennas and subcarriers. Inspired by [16], we parallelize the processing of antennas,i.e., predicting the CSI for each transmitter antenna separately. Thereby, the input sample of $j$-th $\forall j$ antenna can be converted into real tensors and rearranged to merge feature dimensions as $\mathbf{X}_{\rm CP}^{\rm in}\in\mathbb{R}^{T_{1}\times 2M}$. To facilitate network training and convergence, we first normalize the input data as $\mathbf{X}_{\rm CP}^{\rm in^{\prime}}=\frac{\mathbf{X}_{\rm CP}^{\rm in}-\mu_{\rm CP}}{\sigma_{\rm CP}}$, where $(\mu_{\rm CP},\sigma_{\rm CP})$ represent the mean value and standard deviation of a batch of corresponding domain input data. To capture the local temporal features and reduce computational complexity, patching operation along the temporal dimension is adopted. Specifically, $\mathbf{X}_{\rm CP}^{\rm in}$ are grouped into non-overlapping patches of size $N$ as $\mathbf{X}_{\rm CP}^{\rm pat}\in\mathbb{R}^{T^{\prime}_{1}\times N\times 2M}$ [21], where $T^{\prime}_{1}=\lceil\frac{T_{1}}{N}\rceil$ is the number of patches. Zero-padding is applied to the last patch that is not fully filled. We adopt the CSI attention module, which is proposed and described in detail in [22], to extract preliminary temporal and frequency features before LLM. The process can be presented by
@@
where ${\rm CSIATT}^{L}$ represents the CSI attention module cascaded $L$ times. Finally, to adapt to the input format of LLM, $\mathbf{X}_{\rm CP}^{\rm CA}$ is rearranged to $\mathbf{X}_{\rm CP}^{\rm CA^{\prime}}\in\mathbb{R}^{T^{\prime}_{1}\times 2MN}$ and then mapped to the feature dimension of the pre-trained LLM with a single fully-connected layer:
@@
III-C LLM Mainbody with Low-rank Adaptation (LoRA)
The proposed method finetunes LLM by a parameter-efficient technique, low-rank adaptation (LoRA) [23], which freezes the pre-trained model weights and injects trainable rank decomposition matrices into each layer. The use of LoRA mainly brings two advantages. Firstly, due to the immense parameter size of LLMs, full-parameter finetuning could be impractical. LoRA circumvents this by only fine-tuning low-rank adapters, with a significantly smaller number of parameters, yielding results comparable to full-parameter finetuning. Secondly, LoRA prevents catastrophic forgetting of the original knowledge, as the newly learned knowledge has a lower rank than the original knowledge. Therefore, it facilitates us to utilize the universal modeling and generalization capability of pre-trained LLMs to achieve multiple tasks with a single model.
Figure 3: The illustration of LoRA adapters.
Without loss of generality, LLAMA2 is chosen as the LLM backbone in this work. The backbone of LLAMA2 is composed of stacked transformer decoders, the structure of which is illustrated in Fig. 3. We focus on training the linear projection layer and efficient finetuning the LLM backbone using LoRA. With LoRA, we finetune $W_{q}$ and $W_{v}$ via low-rank adaptation. It is worth noting that in the proposed method, the backbone can be flexibly replaced with other LLM, such as GPT [10] or QWen [24]. The selection of the type and size of the LLM, as well as the rank and finetuned modules of LoRA, needs to consider the trade-off between training costs and performance.
Obtaining the LLM backbone, we concatenate the embedding of prompt and data obtained in the previous subsections as $\mathbf{X}^{\rm emb}$ and feed into the LLM backbone. The output can be obtained by
@@ III-D Output Decoders
Similar to input encoders, to facilitate multiple downstream tasks simultaneously, task-specific decoders are required to be designed to convert the output features of the LLM into the final results for different tasks. Then we elaborate on the output decoders of the three tasks sequentially. Note that for output decoders, we discard the prefix portion and retain only the output presentations of the data for the decoder.
@@ III-D1 Decoder for multi-user precoding
As discussed in Section II A., the multi-user precoding problem can be transformed into learning the key features $\mathbf{\lambda}=[\lambda_{1},\lambda_{2},\cdots,\lambda_{K}]$ and $\mathbf{p}=[p_{1},p_{2},\cdots,p_{K}]$. Therefore, a MLP module comprising of two fully-connected (FC) layers is adopted to generates $2K$ values including the power allocation vectors $\hat{\mathbf{\lambda}}$ and $\hat{\mathbf{p}}$. The output of the MLP module $\mathbf{X}^{O}_{\rm Pre}\in\mathbb{R}^{K\times 2}$ can be written as
@@
where $\hat{\mathbf{\lambda}}=\mathbf{X}^{O}_{\rm Pre}[:,1]$ and $\hat{\mathbf{p}}=\mathbf{X}^{O}_{\rm Pre}[:,2]$. Then the scaling layer scales the results of the output layer $\hat{\mathbf{\lambda}}$ and $\hat{\mathbf{p}}$ to meet the power constraint by the following method:
@@ III-D2 Decoder for signal detection
For signal detection, we employ two FC layers to transform the dimension of the $\mathbf{X}^{O}_{\rm LLM}$ to the antenna number:
$\mathbf{X}^{\rm mlp}_{\rm Dec}={\rm MLP}(\mathbf{X}^{O}_{\rm LLM}).$ (23)
Then $\mathbf{X}^{\rm mlp}_{\rm Dec}$ is rearranged $\mathbf{X}^{O}_{\rm Dec}\in\mathbb{R}^{K\times 2}$ to where the first and the second dimension respectively correspond to the real part and the imaginary part.
@@ III-D3 Decoder for channel prediction
For convenience, we also utilize a simple MLP module as the decoder to predict the channel. The output is presented by
@@
Then $\mathbf{X}^{\rm mlp}_{\rm CP}$ is de-normalized to generate the final output of the network, i.e.,
$\mathbf{X}^{norm}_{\rm CP}=\sigma_{\rm CP}\mathbf{X}^{\rm mlp}_{\rm CP}+\mu_{\rm CP}.$ (25)
@@
Then the tensors are rearranged to $\mathbf{X}^{O}_{\rm CP}\in\mathbb{R}^{T_{2}\times M\times 2}$ to separate the real part and the imaginary part of the predicted channel.
III-E Multi-task Loss Function and Training Schedule

### v2

@@ III-B Input Encoders
Adapting text-based pre-trained LLM to multiple communication domain tasks is another challenging problem. To be specific, first, there is a huge gap between the specific characteristics of the data in the communication domain and the natural language. Therefore, directly applying a general-domain LLM to these communication tasks may lead to poor performance. Secondly, the format, length, and meaning of the data also differs significantly among the tasks. The difficulties urge us to design dedicated input encoder modules for each task to fit the feature space of the LLM, and facilitate multi-task feature extraction.
@@ III-B1 Encoder for multi-user precoding
Since neural networks generally deal with real numbers, we first convert the complex multi-user channel into real tensors $\mathbf{X}_{\rm Pre}^{\rm in}\in\mathbb{R}^{K\times 2N_{t}}$ as the input. For multi-user precoding, we employ a transfomer to extract the channel feature and inter-user relationship. The transfomer is composed of $L=3$ blocks. The structure of each block is presented in Fig. 2, including a multi-head self-attention module and a multilayer perceptron (MLP) module. The input $\mathbf{X}_{\rm Pre}^{\rm in}$ is first processed by the multi-head self-attention module:
@@
which serves as the input of next block. Denote the output of the transfomer as $\mathbf{X}_{\rm Pre}^{\rm mlp(3)}$. Then a linear layer is used to project $\mathbf{X}_{\rm Pre}^{\rm mlp(3)}$ into the language model space
@@ III-B2 Encoder for signal detection
The input of signal-detection encoder contains the uplink channel $\bar{\mathbf{H}}^{m}$ and the received signal $\bar{\mathbf{y}}^{m}$. The encoder rearranges the channel $\bar{\mathbf{H}}^{m}$ as one token $\mathbf{X}_{\rm Det}^{\rm Hin}\in\mathbb{R}^{K\times 2N_{t}}$. Then a MLP module is employed to extract features from the channel
$\mathbf{X}_{\rm Det}^{\rm Hmlp}={\rm MLP}(\mathbf{X}_{\rm Det}^{\rm Hin}).$ (17)
We also extract shallow features from the rearranged real received signal $\mathbf{X}_{\rm Det}^{\rm yin}$ by a MLP module
$\mathbf{X}_{\rm Det}^{\rm ymlp}={\rm MLP}(\mathbf{X}_{\rm Det}^{\rm yin}).$ (18)
@@
$\mathbf{X}_{\rm Det}^{\rm emb}={\rm Linear}([\mathbf{X}_{\rm Det}^{\rm Hmlp},\mathbf{X}_{\rm Det}^{\rm ymlp}]).$ (19)
@@ III-B3 Encoder for channel prediction
Consider that the CSI $\mathbf{H}^{t}\in\mathbb{C}^{N_{t}\times M}$ of at time $t$ is the high-dimensional structural data, directly predicting the matrix by the network will bring significant complexity. The complexity can be unacceptable for future 6G systems with a large number of antennas and subcarriers. Inspired by [13], we parallelize the channel prediction for different antennas. That is to say, we predict the CSI of each transmitter antenna separately. Thereby, the input sample of $j$-th antenna, for $j\in\{1,2,\cdots,N_{T}\}$, can be converted into a real tensor $\mathbf{X}_{\rm CP}^{\rm in}\in\mathbb{R}^{T_{1}\times 2M}$. To facilitate convergence of network training, we first perform batch normalization for the input data as $\mathbf{X}_{\rm CP}^{\rm in^{\prime}}$. Then the normalized input $\mathbf{X}_{\rm CP}^{\rm in^{\prime}}$ is divided into $T^{\prime}_{1}=\lceil\frac{T_{1}}{N}\rceil$ non-overlapping patches of size $N$ along the temporal dimension [24]. The patching operation helps to capture the local temporal features and the patched input is denoted as $\mathbf{X}_{\rm CP}^{\rm pat}\in\mathbb{R}^{T^{\prime}_{1}\times N\times 2M}$. We adopt the CSI attention module proposed in [25], to extract preliminary temporal and frequency features before LLM:
@@
where ${\rm CSIATT}^{L}$ represents the CSI attention module cascaded $L$ times. Finally, to align with the input format of LLM backbone, $\mathbf{X}_{\rm CP}^{\rm CA}$ is rearranged to $\mathbf{X}_{\rm CP}^{\rm CA^{\prime}}\in\mathbb{R}^{T^{\prime}_{1}\times 2MN}$ and then mapped to the feature dimension of the pre-trained LLM with a single fully-connected layer:
@@
III-C LLM Mainbody
Figure 3: The illustration of a transformer block in LLAMA2.
In this paper, we adopt LLAMA2 as the LLM backbone for our proposed multi-task PHY framework. It should be emphasized that our framework allows for seamless integration of alternative LLMs, including but not limited to GPT [10] and QWen [26]. The decision of model architecture and scale requires evaluation of the trade-off between computational complexity and performance.
The architecture of LLAMA2 is presented in Fig. 3. It consists of stacked classical transformer decoders. Note that in the proposed multi-task PHY network, the LLM backbone takes up most of the model parameters. In contrast, the encoders and decoders, though designed specifically for each task, only occupies a very small portion in the network. To be specific, the parameter size of LLAMA2-7B backbone is 7 billion, while the parameter size of all other modules, including the multi-task instructions, input encoders and output decoders, is only about 19 million. Obtaining the LLM backbone, we concatenate the embedding of prompt and data obtained in the previous subsections as $\mathbf{X}^{\rm emb}$ and feed into the LLM backbone. The output can be obtained by
@@ III-D Output Decoders
Similar to input encoders, to facilitate multiple downstream tasks simultaneously, task-specific decoders are required to convert the output features of the LLM into the final results for different tasks. Here, we elaborate on the output decoders of the three tasks sequentially. Note that for output decoders, we discard the prefix portion that is associated with the instruction prompt, while we only retain the output presentations of the data for the decoder.
@@ III-D1 Decoder for multi-user precoding
As discussed in Section II A., the multi-user precoding problem can be transformed into the learning of the parameters $\mathbf{\lambda}=[\lambda_{1},\lambda_{2},\cdots,\lambda_{K}]$ and $\mathbf{p}=[p_{1},p_{2},\cdots,p_{K}]$. Therefore, a MLP module comprising of two fully-connected (FC) layers is adopted to generates $2K$ values including the power allocation vectors $\hat{\mathbf{\lambda}}$ and $\hat{\mathbf{p}}$. The output of the MLP module $\mathbf{X}^{O}_{\rm Pre}\in\mathbb{R}^{K\times 2}$ can be written as
@@
where $\hat{\mathbf{\lambda}}=\mathbf{X}^{O}_{\rm Pre}[:,1]$ and $\hat{\mathbf{p}}=\mathbf{X}^{O}_{\rm Pre}[:,2]$. Then the scaling layer scales the results of the output layer $\hat{\mathbf{\lambda}}$ and $\hat{\mathbf{p}}$ to meet the power constraint by:
@@ III-D2 Decoder for signal detection
For signal detection, we employ two FC layers to transform the dimension of the $\mathbf{X}^{O}_{\rm LLM}$ to the number of antennas:
$\mathbf{X}^{\rm mlp}_{\rm Det}={\rm MLP}(\mathbf{X}^{O}_{\rm LLM}).$ (25)
Then $\mathbf{X}^{\rm mlp}_{\rm Det}$ is rearranged $\mathbf{X}^{O}_{\rm Det}\in\mathbb{R}^{K\times 2}$ to where the first and the second dimension respectively correspond to the real part and the imaginary part.
@@ III-D3 Decoder for channel prediction
We utilize an MLP module as the decoder to predict the channel. The output is presented by
@@
Last, $\mathbf{X}^{\rm mlp}_{\rm CP}$ is de-normalized to generate the final output of the network, i.e.,
$\mathbf{X}^{norm}_{\rm CP}=\sigma_{\rm CP}\mathbf{X}^{\rm mlp}_{\rm CP}+\mu_{\rm CP},$ (27)
@@
where $\mu_{\rm CP}$ and $\sigma_{\rm CP}$ is the mean and variance of the channel matrix. Then the tensors are rearranged to $\mathbf{X}^{O}_{\rm CP}\in\mathbb{R}^{T_{2}\times M\times 2}$ to separate the real part and the imaginary part of the predicted channel.
IV The Fine-tuning strategy of the Proposed Physical Layer Multi-task LLM
Considering the huge parameter size of LLMs, directly fine-tuning all the parameters of the proposed network is impractical. To address this issue, this section focuses on the computationally efficient fine-tuning of the proposed PHY multi-task LLM. First, we introduce LoRA [20] for LLM fine-tuning, which inserts low-rank adapters into the LLM backbone to finetune the pre-trained LLM for PHY tasks. Then, to mitigate the computational and memory demands of the proposed model, we employ a LoRA fine-tuning-aware quantization method [21] that performs LLM quantization while concurrently properly initializes the low-rank adapters for LoRA-based fine-tuning. Moreover, the multi-task loss as well as training schedule are illustrated.
IV-A LoRA for LLM Mainbody
LoRA [20] is a parameter-efficient technique, which freezes the pre-trained model weights and injects trainable rank decomposition matrices into each layer. The use of LoRA mainly brings two advantages. Firstly, due to the immense parameter size of LLMs, full-parameter fine-tuning could be impractical. LoRA circumvents this by only fine-tuning low-rank weight matrices, with a significantly reduced number of parameters. Secondly, LoRA prevents the problem of catastrophic forgetting of the original knowledge during fine-tuning. This is attributed to the fact that the newly learned knowledge has a lower rank than the original knowledge. As a result, LoRA facilitates the use of the universal modeling and generalization capability of pre-trained LLMs to achieve multiple tasks with a single model.
Specifically, we focus on the fine-tuning of the query matrices and value matrices in an attention block of the LLM backbone, denoted as $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ respectively, as illustrated in Fig. 3. LoRA updates two rank decomposition weight matrices $\mathbf{A}$ and $\mathbf{B}$ that are attached to a frozen pre-trained weight matrix $\mathbf{W}$ ($\mathbf{W}_{q}$ and $\mathbf{W}_{v}$). In this case, the original weight $\mathbf{W}$ is modified as
$\mathbf{W}\leftarrow\mathbf{W}+\mathbf{A}\mathbf{B}^{T},$ (28)
where $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$, $\mathbf{A}\in\mathbb{R}^{d_{1}\times r}$, $\mathbf{B}\in\mathbb{R}^{d_{2}\times r}$, and $r\ll\min\{d_{1},d_{2}\}$. Generally, we initialize the weights as follows:
$\mathbf{A}\sim\mathcal{N}(0,\sigma^{2}),\quad\mathbf{B}=0.$ (29)
As mentioned above, during the fine-tuning, we freeze $\mathbf{W}$ while update $\mathbf{A}$ and $\mathbf{B}$. Since $r<<\min(d_{1},d_{2})$, the number of parameters for fine-tuning $\mathbf{A}$ and $\mathbf{B}$, i.e., $rd_{1}+rd_{2}$, is significantly less than that of the full weight matrix, $d_{1}d_{2}$. The reduced parameter size thus makes the LoRA-based fine-tuning much efficient.
IV-B LoRA Fine-tuning-aware Quantization
Despite the usage of LoRA, the extensive computational and memory demands of LLM-based models still pose significant challenges for both fine-tuning and inference, especially for resource-restricted equipment in wireless communications. To reduce the storage demands of pre-trained models, quantization acts as an essential compression strategy. Its key idea is to transform the original weights with a high precision into a finite set of discrete values. For instance, converting model parameters from the original 16-bit floating-point format (FP16) to a 4-bit integer format leads to a 75% decrease in storage overhead. However, it is worth noting that a quantized network parameters might degrade the performance of the aforementioned LoRA fine-tuning strategy. Particularly, consider a quantized weight matrix $\mathbf{Q}=q(\mathbf{W})$, where $q(\cdot)$ denotes the quantization operator. If the low-rank adapters $\mathbf{A}$ and $\mathbf{B}$ initialized by (29) and then attached to $\mathbf{Q}$, the initial weight matrix $\mathbf{Q}+\mathbf{A}\mathbf{B}^{T}$ naturally diverges from the original pre-trained parameters $\mathbf{W}$ because of the variations caused by the quantization process. Such unavoidable variations may negatively influence the initial setup of the LoRA fine-tuning procedure.
To solve the problem, we adopt a LoRA-fine-tuning-aware method to smoothly integrate quantization into the procedure of LoRA fine-tuning. The main idea is to approximate the original high-precision pre-trained weights by alternatively applying quantization for LLMs and proper low-rank initialization for LoRA fine-tuning. This initialization strategy effectively reduces the gap between the quantized and full-precision model, leading to a notable enhancement in fine-tuning performance.
Figure 4: The comparison of LoRA fine-tuning-aware quantization with traditional full-precision LoRA.
The key procedures of the LoRA fine-tuning-aware quantization method are presented in Fig. 4. During the initialization of LoRA fine-tuning, a quantized weight matrix $\mathbf{Q}\in\mathbb{R}^{d_{1}\times d_{2}}$ with $N$-bit precision, along with the low-rank adapters $\mathbf{A}\in\mathbb{R}^{d_{1}\times r}$, $\mathbf{B}\in\mathbb{R}^{d_{2}\times r}$ are designed to closely approximate the original full-precision pre-trained parameter matrix $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$. Mathematically, the model weight initialization problem can be formulated as:
[…]
This problem can be efficiently solved via alternatively conducting quantization and singular-value-decomposition. The step-by-step procedures provided in Algorithm 1 and the details are as follows.
Quantization Step: In the $i$-th iteration, the quantization process is applied to the residual between the pre-trained parameter matrix $\mathbf{W}$ and the low-rank approximation $\mathbf{A}_{i-1}\mathbf{B}_{i-1}^{T}$, yielding the quantized weight matrix $\mathbf{Q}_{i}$
$\mathbf{Q}_{i}=q_{N}(\mathbf{W}-\mathbf{A}_{i-1}\mathbf{B}_{i-1}^{T}),$ (31)
where $q_{N}(\cdot):\mathbb{R}\mapsto\{0,1,\cdots,2^{N}-1\}$ maps a high-precision weight matrix, e.g., matrix with 16-bit floating point number, to an $N$-bit quantized matrix. Typically, the quantization process can be expressed as
$\mathbf{Q}={\rm round}((2^{N}-1)F(\mathbf{W})),$ (32)
where $F(\cdot):\mathbb{R}\mapsto[0,1]$ is a normalization function. In this work, we utilize the 4-bit NormalFloat Quantization (NF4) proposed in [27] to model the normalization function. It assumes $\mathbf{W}\sim N(0,\sigma^{2})$ and hence $F(\mathbf{W})=\Phi(\mathbf{W}/\sigma)$, where $\Phi(\cdot)$ is the cumulative distribution function of the standard normal distribution. Besides, other quantization methods can also be involved such as the uniform quantization.
SVD Step: After obtaining the $i$-th quantized weight $\mathbf{Q}_{i}$, SVD is applied to the residual of the quantization denoted by $\mathbf{R}_{i}=\mathbf{W}-\mathbf{Q}_{i}$ by
$\mathbf{R}_{i}=\sum_{j=1}^{d}\sigma_{i,j}{\bf u}_{i,j}{\bf v}_{i,j}^{T},$ (33)
where $d=\min\{d_{1},d_{2}\},\quad\sigma_{i,1}\geq\sigma_{i,2}\geq\ldots\geq\sigma_{i,d}$ are the singular values of $\mathbf{R}_{i}$, ${\bf u}_{i,j}$, ${\bf v}_{i,j}$ are the associated left and right singular vectors of $\mathbf{R}_{i}$. We then obtain a rank-$r$ approximation of $\mathbf{R}_{i}$ by $\mathbf{A}_{i-1}\mathbf{B}_{i-1}^{T}$, where
$\mathbf{A}_{i}=[\sqrt{\sigma_{i,1}}{\bf u}_{i,1},\ldots,\sqrt{\sigma_{i,r}}{\bf u}_{i,r}],$ (34)
$\mathbf{B}_{i}=[\sqrt{\sigma_{i,1}}{\bf v}_{i,1},\ldots,\sqrt{\sigma_{i,r}}{\bf v}_{i,r}].$ (35)
After obtaining the initialization, LoRA fine-tuning can be performed as described in Section III-C.
Moreover, during the inference, we merge the quantized backbone with the finetuned adapters to acquire the final output. If the deployed device is resource-limited, the merged model can be further quantized before deployment; otherwise, the full-precision model can be deployed directly.
Algorithm 1 The initialization of LoRA fine-tuning-aware quantization.
0:  Full-precision pre-trained weight $\mathbf{W}$, LoRA rank $r$, $N$-bit quantization function $q_{N}(\cdot)$, iteration number $Iter$.
1:  Initialize $\mathbf{A}_{0}\leftarrow\mathbf{0},\mathbf{B}_{0}\leftarrow\mathbf{0}$;
2:  for $i=1$ to $Iter$ do
3:   Obtain quantized weight $\mathbf{Q}_{i}\leftarrow q_{N}(\mathbf{W}-\mathbf{A}_{i-1}\mathbf{B}_{i-1}^{T})$;
4:   Obtain low-rank approximation $\mathbf{A}_{i},\mathbf{B}_{i}\leftarrow$ SVD$(\mathbf{W}-\mathbf{Q}_{i})$ based on [36]-[37];
5:  end for
5:  $\mathbf{Q}_{Iter}$, $\mathbf{A}_{Iter}$, and $\mathbf{B}_{Iter}$.
IV-C Multi-task Loss Function and Training Schedule